OpenStudy (anonymous):

Find f(5) for this sequence: (2 points) f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ please helpppp

OpenStudy (mayankdevnani):

find f(3) first ?

OpenStudy (mayankdevnani):

can you find it ? @mathing64

OpenStudy (anonymous):

does f(3) = 12?

OpenStudy (mayankdevnani):

great !

OpenStudy (mayankdevnani):

and what about f(4) ?

OpenStudy (mayankdevnani):

@mathing64

OpenStudy (anonymous):

does f(4) = 19??

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

great !

OpenStudy (mayankdevnani):

now what about f(5) ? @mathing64

OpenStudy (mayankdevnani):

i know you can find out ?

OpenStudy (anonymous):

37?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

nope !

OpenStudy (anonymous):

what would it be then?

OpenStudy (mayankdevnani):

\[\large \bf f(5)=f(1)+f(2)+f(5-1)=f(1)+f(2)+f(4)\] \[\large \bf f(5)=2+5+19=26\]

OpenStudy (mayankdevnani):

understood ?

OpenStudy (anonymous):

ohhhh i see where i went wrong now thanks so much for all the help!

OpenStudy (mayankdevnani):

your welcome :)